Two trains(both are 400m in lenght) are travelling on two parallel tracks with the same speed 72kmph.train A is ahead of train B. Now train B driver decides to overtake train A and accelarates at 1m/s2


and manages to overtake in 50 sec.Now calculate the initial seperation between the trains before the train B's accelaration?


 

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear karthik

let the initial seperation = x

In train A frame of reference .velocity of train A is zero.initia velocity of train B is zero.

to overtake train A ,train A has to cover L+800 distance.

so

 S =1/2at2

L+800 =1/2 *1 * (50)2

L=450 m

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin